a rectangle is three times as long as it is wide. write and simplify an expression for the perimeter of the rectangle in the terms of the width "w".
Answer by edjones(7311)

L=3w
2L+2w=P
2(3w)+2w=P substitute 3w for L.
6w+2w=P
8w=P
.
Ed